OpenStudy (star):

find the particular solution of (x^2 + y^2)y' + 2xy = 0 satisfying y(1)=1

OpenStudy (star):

\[\left( x ^{2} + y^{2} \right) y \prime +2xy = 0\]

OpenStudy (anonymous):

it is a homogenous differential equation \[dy/dx=\-\2xy/x \^\{\2\}\+y \^\{\2\}\]

OpenStudy (anonymous):

subs y=vx

OpenStudy (anonymous):

the answer is y=lnx+(1/3)ln((y/x)^3 +3(y/x))+C y(1)=1 1=ln1+(1/3)ln(1^3 +3)+C 1=0+(1/3)ln4 +C C=1-(1/3)ln4=0.538 y=lnx+(1/3)ln((y/x)^3 +3(y/x))+0.538